OpenStudy (anonymous):

c+(4-3c)-2=0 help me solve this plez

OpenStudy (anonymous):

\[c + (4-3c)-2 = 0\] parenthesis unnecessary: \[c + 4 - 3c - 2 = 0\] combine like terms: \[-2c + 2 = 0\] isolate c: \[-2c = -2\] divide by -2: \[c = 1\]

OpenStudy (anonymous):

how u get the -2c+2=0